# Optimizing the Network Topology of a Linear Reservoir Computer

### v1

@@ 2.2 Optimization
Having established the relationships between the errors and their formulations, we now propose an optimization framework to minimize the error of the linear reservoir computer by identifying an optimal set of eigenvalues $\lambda_{i}$ in Eqs. (7) and (8). The primary goal of this procedure is to find these eigenvalues along with the parameter $\kappa$ that jointly minimizes both the training and testing errors.
@@
$H=\frac{N}{\sum_{\begin{subarray}{c}j,z=1\\ j\neq z\end{subarray}}^{N}\frac{1}{|\lambda_{j}-\lambda_{z}|}}$ (30h)
@@
In the optimization, it is desired that the eigenvalues be distinct, as this prevents dynamic degeneracy, allowing the reservoir to generate richer and more diverse responses to inputs. The cost function Eq. (30a) includes three terms. The first term represents the reservoir’s error in the frequency domain and serves as the primary quantity to be minimized. High output readouts often indicate overfitting and can amplify noise in the reservoir states or inputs, leading to poor generalization and reduced robustness. To mitigate this, a second term is included in the cost function that minimizes the norm of the output readouts, where $\boldsymbol{\kappa}$ is the vector of output readout weights, and $\beta_{1}$ is a positive constant controlling the weight of this penalty. The third term is the reciprocal of the harmonic mean of the differences between the eigenvalues, which is lower when the difference in eigenvalues is high. It is included in the cost function to enforce distinctness among the eigenvalues, and $\beta_{2}$ is the weight that controls the strength of this term.
@@ 3 Numerical Verifications
We validate the theoretical results through numerical simulations. Reservoir computer simulations are carried out in both MATLAB and Julia. The optimization problem is numerically solved using Ipopt, a nonlinear optimization solver, in combination with the JuMP modeling framework in Julia, following Algorithm 1. The procedure in the Algorithm 1 is repeated with different initial values for $\bm{\lambda}_{0}$, which have been chosen from a uniform distribution over the interval [20,0][-20,0] , and the best performance is stored as the optimal solution. The vector $\bm{\kappa}_{0}$ is then obtained using ridge regression, while $\bm{M}_{0}$, and $\bm{\theta}_{0}$ are computed according to Eq. (12). (The implementation code is publicly available at here)
@@
[…]
7:for $iter=1$ to num_iternum\_iter do
8:  Cost = $\min\sum_{i=1}^{K}W_{i}(\epsilon_{\cos,i}^{2}+\epsilon_{\sin,i}^{2})+\beta_{1}\sum_{j=1}^{N}\kappa_{j}^{2}+\beta_{2}(1/H)$
9:  Add nonlinear constraints
10:  Set initial values for $\lambda,\kappa,M,\theta$ from the previous iteration or initial guess.
11:  Solve the optimization problem and find $\bm{\lambda}_{new},\bm{\kappa}_{new},\bm{M}_{new},\bm{\theta}_{new}$
12:  if solver converged successfully then
13:   $\epsilon=\sqrt{\sum_{i}W_{i}(\epsilon_{\cos,i}^{2}+\epsilon_{\sin,i}^{2})}$
14:   if $\epsilon<lowest\_error$ then
15:     $lowest\_error\leftarrow\epsilon$
16:     Store current $\bm{\lambda}_{new},\bm{\kappa}_{new},\bm{M}_{new},\bm{\theta}_{new}$ as the best solution.
17:   end if
18:  else
19:   Keep previous $\bm{\lambda}_{best},\bm{\kappa}_{best},\bm{M}_{best},\bm{\theta}_{best}$ as the current solution.
20:  end if
21:end for
22:return $\bm{\lambda}_{best},\bm{\kappa}_{best},\bm{M}_{best},\bm{\theta}_{best},lowest\_error$
@@ 3.1 Optimal Reservoir Computer
To demonstrate the effectiveness of the proposed optimization framework, two numerical examples are considered involving a reservoir computer governed by Eq. (3): one with $N=10$ nodes and the other with $N=100$ nodes. In both examples, the reservoir is trained for 3000 steps with parameters $\beta_{1}=10^{-7}$, $\beta_{2}=10^{-1}$, and $\gamma=6$, and the reservoir states together with the corresponding errors are computed using both analytical and numerical methods. The matrix $A$ is taken to be the adjacency matrix of a weighted/unweighted Erdos-Renyi graph with its spectrum shifted to ensure the Hurwitz property, and the input mask vector $c$ has entries randomly drawn from a normal distribution. The input and output signals are defined as follows:
@@
Figure 2: Performance of the 10-node reservoir. Red dashed lines represent the input signal, while blue lines correspond to the reservoir computer output: (a) Training before optimization, (b) Testing before optimization, (c) Training after optimization, (d) Testing after optimization.
The optimization formulation is also effective for larger networks. Figure (3) illustrates the reservoir’s performance before and after the optimization for a $N=100$-node network.
Figure 3: Performance of the 100-node reservoir. Red dashed lines represent the input signal, while blue lines correspond to the reservoir computer output: (a) Training before optimization, (b) Testing before optimization, (c) Training after optimization, (d) Testing after optimization.
According to the Figures (2) and (3), the proposed optimization shows significant potential for improving signal estimation accuracy.
@@ 3.2 Performance Relative to Existing Approaches
To evaluate the performance of the proposed optimal LRC, it is compared with a benchmark nonlinear reservoir computer reported in Gauthier et al. (2021), using two activation functions: “tanh” and “ReLU”, as well as with a linear reservoir computer benchmark. Fig. (4) presents the training NRMSE of the optimized reservoir alongside the three benchmarks under two scenarios. In the first scenario, the number of reservoir nodes is fixed at 50 while the number of input frequencies varies; in the second, the number of input frequencies is fixed at 3 while the number of reservoir nodes varies. These cases allow assessment of the impact of network size and input complexity on performance and highlight the advantages of the proposed optimal reservoirs over the benchmarks.
@@
Figure 4: Average normalized root mean squared error (NRMSE) of reservoir computers with their corresponding standard deviations. The yellow line represents the NRMSE of the linear benchmark, the orange line corresponds to the NLRC with ReLU activation, the blue line shows the NLRC with tanh activation, and the purple line indicates the optimized reservoir computer. a) The number of nodes is fixed at 50 while the number of input frequencies increases. b) The number of input frequencies is fixed at 3 while the number of reservoir nodes increases.
Figure (4)(a) shows that the optimized LRC achieves the lowest training error, while the linear benchmark exhibits the highest. The NLRC with tanh activation performs better than the optimized LRC when the number of input frequencies is small; however, its error increases sharply as the frequency count grows, eventually exceeding that of the optimized LRC. In contrast, the optimized LRC shows only a modest increase in error under the same conditions, demonstrating greater robustness to input complexity and stronger generalization than the other methods. The NLRC with ReLU activation performs worse than the tanh-based reservoir but still outperforms the linear benchmark.
Figure (4)(b) shows that the optimized LRC achieves the lowest NRMSE when the reservoir size is small. As the number of nodes increases, however, the NRMSE of the tanh-based NLRC decreases drastically, eventually surpassing the optimized LRC, while the linear benchmark maintains the highest NRMSE with only minor improvements. The ReLU-based NLRC exhibits consistently higher errors than both the optimized LRC and the tanh-based NLRC, with only gradual reduction as the number of nodes increases. Importantly, the optimized LRC achieves significantly lower NRMSE with far fewer nodes compared to the tanh-based NLRC. This reduction in required network size directly translates into lower computational cost, making the proposed optimized LRC more efficient for real-time and resource-constrained applications.
@@ 3.3 Effect of the Cost Function Components on the Optimization Performance
The optimization formulation, as shown in Eq. (30), consists of three terms. The second and third terms have associated weights $\beta_{1}$ and $\beta_{2}$, respectively. Therefore, it is important to analyze the impact of these coefficients on the overall optimization process. Figure (5) (a) demonstrates the effect of the second and third terms’ weights on the optimization of the error.
Figure (5) (a), shows that reducing the weight of the second term ($\beta_{1}$) leads to a decrease in optimization error. This suggests that allowing larger output weights enables the reservoir to more effectively learn the signal, thereby improving performance. Additionally, increasing the weight of the third term ($\beta_{2}$) also leads to a reduction in error. This is because the third term encourages the optimization to produce distinct eigenvalues, which is a desirable property.
@@ 3.4 Effect of $\lambda$ on Reservoir Performance
Performing a sensitivity analysis of the reservoir’s eigenvalues is important for assessing both robustness and generalization. High sensitivity means that small perturbations in eigenvalues can lead to large errors, making the system fragile. It may also indicate overfitting, where the reservoir performs well only under narrowly tuned conditions. Evaluating sensitivity helps ensure stable and reliable performance. Since the optimization formulation focuses on minimizing the error by tuning the reservoir’s eigenvalues and output weights, it is essential to understand how sensitive the reservoir computer is to the selection of these optimal eigenvalues. In order to do the sensitivity analysis, a small perturbation according to Eq.(31) is added to the eigenvalues.
$\lambda_{perturbed}=\lambda_{opt}+\epsilon_{s}\delta,\quad\delta\in(-1,0)$ (31)
Figure 5: a) Effect of optimization terms’ weights on the optimization error. b) Sensitivity of the Reservoir Training Error to Optimal Eigenvalues.
Figure (5) (b) indicates that the reservoir computer’s learning error remains largely unaffected by small perturbations in the optimal eigenvalues. Only significant deviations from these optimal values cause noticeable increases in the error. This behavior suggests that the reservoir is robust and stable around the optimized eigenvalues, making it resilient to minor noise or implementation inaccuracies. Consequently, the optimized reservoir parameters provide a reliable solution, ensuring consistent performance even when slight variations occur in practice.
@@ 4 Conclusion
There are two main contributions of this work. First, we demonstrate that the dynamics of a linear RC can be decoupled into a a number of independent modes while preserving a comparable error to the coupled RC. Second, we show that each of these modes can be individually optimized to perform a given task of interest when the decoupled RC is reformulated from the time domain into the frequency domain. This reformulation allows the reservoir modes matrix to be significantly reduced, since the number of input frequencies is typically much smaller than the number of time steps. We show that the effect of the connectivity of the reservoir on the training and testing error is completely captured by the eigenvalues of the reservoir adjacency matrix, which allows us to formulate an optimization problem in the eigenvalues only. Simulations on networks of varying sizes demonstrate that the optimized RC significantly outperforms randomly constructed reservoirs in both the training and testing phases. Moreover, comparisons with established reservoir computing benchmarks indicate that the optimized reservoir not only achieves superior performance but also exhibits greater robustness to input complexity, stronger generalization capabilities, and higher efficiency for real-time and resource-constrained applications.

### v2

@@ 2.2 Optimization
Having established the relationships between the errors (22), we now propose an optimization framework to minimize the error of the linear reservoir computer by identifying an optimal set of eigenvalues $\lambda_{i}$ in Eqs. (7) and (8). The primary goal of this procedure is to find these eigenvalues along with the output weights $\boldsymbol{\kappa}$ that jointly minimizes the training error, as we explain below.
@@
$H=\frac{N}{\sum_{\begin{subarray}{c}j,z=1\\ j\neq z\end{subarray}}^{N}\frac{1}{|\lambda_{j}-\lambda_{z}|}},$ (25h)
@@
The cost function Eq. (25a) includes three terms. The first term represents the reservoir’s error in the frequency domain and serves as the primary quantity to be minimized. High output readouts often indicate overfitting and can amplify noise in the reservoir states or inputs, leading to poor generalization and reduced robustness. To mitigate this, a second term is included in the cost function that minimizes the norm of the output readouts, where $\boldsymbol{\kappa}$ is the vector of output readout weights, and $\beta_{1}\geq 0$ is a constant controlling the weight of this penalty.
Some comments should be made regarding the third term. The third term is the reciprocal of the harmonic mean of the differences between the eigenvalues, which is lower when any of the differences in eigenvalues are high. It is included in the cost function to enforce distinctness among the eigenvalues, and $\beta_{2}\geq 0$ is the weight that controls the strength of this term. Our analysis shows that identical eigenvalues would produce a computational redundancy and reduce the effective dimension of the reservoir. It was found that without the heuristic term, sometimes a numerical issue would create eigenvalue copies. Considering that this is undesirable, a heuristic algebraic barrier was included to keep the eigenvalues from being identical. The harmonic mean is a simple algebraic way of doing this because of the denominator of the differences of eigenvalues. When the eigenvalues are not near each other, the harmonic term does not contribute significantly and when the eigenvalues come close in value, the harmonic term becomes a significant contribution to the objective function. It is equivalent to a barrier method.
@@ 3 Numerical Verifications
We validate the theoretical results through numerical simulations. Reservoir computer simulations are carried out in both MATLAB and Julia. The optimization problem is numerically solved using Ipopt, a nonlinear optimization solver, in combination with the JuMP modeling framework in Julia, following Algorithm 1, and the best performance is stored as the optimal solution. The optimization procedure in Algorithm 1 is nonconvex; therefore, it is executed multiple times with different initial values of $\lambda_{0}$. As the number of reservoir nodes increases, the dimensionality of the optimization problem also grows, which makes the search more challenging and necessitates a larger number of initial points to reliably approach an optimal solution.
Two strategies are used in parallel to generate the initial guesses. The first strategy is a warm-start approach, in which Algorithm 1 is repeatedly executed using various initial values of the eigenvalues $\bm{\lambda}_{0}$, which we draw from a uniform distribution in the interval [20,0][-20,0]. For each selected $\lambda_{0}$, the optimization is run once, and the resulting optimized value is then used as the initial guess for the next iteration to escape local minima. This process continues until the number of initial points drawn from the uniform distribution reaches the predefined maximum number for the warm-start optimization.
The vector $\bm{\kappa}_{0}$ is then obtained using ridge regression, while $\bm{M}_{0}$, and $\bm{\theta}_{0}$ are computed according to Eq. (15). (The implementation code is made publicly available and can be downloaded here)
@@
6:for $iter=1$ to num_iternum\_iter do
7:  Cost = $\min\sum_{i=1}^{K}W_{i}\epsilon_{q,i}^{2}+\beta_{1}\sum_{j=1}^{N}\kappa_{j}^{2}+\beta_{2}(1/H)$
8:  Add nonlinear constraints
9:  Set initial values for $\lambda,\kappa,M,\theta$ from the previous iteration or initial guess.
10:  Solve the optimization problem and find $\bm{\lambda}_{new},\bm{\kappa}_{new},\bm{M}_{new},\bm{\theta}_{new}$
11:  if solver converged successfully then
12:   $\epsilon=\sqrt{\sum_{i}W_{i}\epsilon_{q,i}^{2}}$
13:   if $\epsilon<lowest\_error$ then
14:     $lowest\_error\leftarrow\epsilon$
15:     Store current $\bm{\lambda}_{new},\bm{\kappa}_{new},\bm{M}_{new},\bm{\theta}_{new}$ as the best solution.
16:   end if
17:  else
18:   Keep previous $\bm{\lambda}_{best},\bm{\kappa}_{best},\bm{M}_{best},\bm{\theta}_{best}$ as the current solution.
19:  end if
20:end for
21:return $\bm{\lambda}_{best},\bm{\kappa}_{best},\bm{M}_{best},\bm{\theta}_{best},lowest\_error$
@@ 3.1 Optimal Reservoir Computer
To demonstrate the effectiveness of the proposed optimization framework, three numerical examples are considered for a reservoir computer governed by Eq. (3), with number of nodes $N=10$, $N=100$, and $N=200$. In the three examples, the reservoir is trained for $T=3000$ steps with parameters $\beta_{1}=10^{-7}$, $\beta_{2}=10^{-1}$, and $\gamma=6$, and the reservoir states together with the corresponding errors are computed using both analytical and numerical methods. The matrix $A$ is taken to be the adjacency matrix of a weighted/unweighted Erdos-Renyi graph with its spectrum shifted to ensure the Hurwitz property, and the input mask vector c has entries randomly drawn from a Normal distribution. We set a larger number of warm-start guesses as the reservoir size grows, namely the warm-start count is set to 50 for the $N=10$ reservoir, 400 for $N=100$, and 1500 for $N=200$. In all cases, the length of each warm-start chain is fixed at 40. Consequently, the total number of initial guesses is equal to $50\times 40=2000$ for $N=10$, $400\times 40=16{,}000$ for $N=100$, and $1400\times 40=60{,}000$ for $N=200$. The input and output signals are defined as follows:
@@
Figure 2: Performance of the 10-node reservoir. Red dashed curves represent the training signal, while blue curves correspond to the best fit produced by the reservoir: (a) Training before optimization, (b) Testing before optimization, (c) Training after optimization, (d) Testing after optimization.
The optimization method also scales to larger networks. Figure (3) shows that, for a $N=100$ reservoir, the unoptimized network fails to learn the signal and exhibits large errors, while the optimized network accurately captures the dynamics and reproduces the target output.
Figure 3: Performance of the 100-node reservoir. Red dashed curves represent the training signal, while blue curves correspond to the best fit produced by the reservoir: (a) Training before optimization, (b) Testing before optimization, (c) Training after optimization, (d) Testing after optimization.
Figures (2) and (3) seem to indicate that the proposed optimization has significant potential for improving signal estimation accuracy.
Average NRMSE
Network Nodes Training Before Optimization Training After Optimization Testing Before Optimization Testing After Optimization
10 0.7849 0.0058 0.6943 0.0054
100 0.5786 0.0053 0.5197 0.0049
200 0.5415 0.0049 0.4983 0.0047
Table 1: Average training and testing NRMSE before and after optimization for different network sizes. The NRMSE values presented in the Table are averaged over 50 independent runs.
The comparison of the average training and testing NRMSE of reservoir computers with 10, 100, and 200 nodes over 50 runs is presented in Table (1). The results show that the optimized reservoirs consistently achieve substantially lower NRMSE across all network sizes. Moreover, increasing the number of nodes does not degrade performance; instead, the optimized reservoirs maintain strong accuracy as the number of nodes increases.
The proposed optimal RC performs reliably not only on simple benchmark signals but also on signals derived from a nonlinear system. Figure (4) demonstrates its performance in an observation task in which the $z$-state of a chaotic Lorenz oscillator is reconstructed from knowledge of its $x$-state (simulation parameters are provided in the Supplement). First, the $K=6$ dominant frequencies common to the $x$ and $z$ states are extracted using the method described in the section S4 of the supplement. The reservoir is then trained to estimate $z$ using only the observed $x$ data.
For this task, the non-optimal RC achieves a training NRMSE of $0.9549$, whereas the optimal RC reduces this error to $0.0141$. The test NRMSE follows the same trend: $0.9521$ for the non-optimal RC and $0.0127$ for the optimal RC. These results clearly highlight the significant performance advantage of the optimal RC.
Figure 4: Performance of an $N=25$-node reservoir used to observe the dynamics of the chaotic Lorenz system. Both the input and training signals are restricted to the $K=6$ dominant common frequencies. Red dashed curves represent the training signal, while blue curves correspond to the best fit produced by the reservoir: (a) Training before optimization, (b) Testing before optimization, (c) Training after optimization, (d) Testing after optimization.
@@ 3.2 Performance Relative to Existing Approaches
To evaluate the performance of the proposed optimal LRC, it is compared with two benchmark nonlinear reservoir computers reported in the supplementary material, using two activation functions: “tanh” and “ReLU”, as well as with a linear reservoir computer benchmark. Fig. (5) presents the testing NRMSE of the optimized reservoir alongside the three benchmarks, under two scenarios. In the first scenario, the number of reservoir nodes is fixed at $N=50$, while the number $K$ of input frequencies varies; in the second, the number of input frequencies is fixed at $K=10$, while the number of reservoir nodes $N$ varies. Together, these scenarios provide a structured basis for evaluating the relative performance of the proposed optimal LRC across changes in the network size and input complexity.
@@
Figure 5: The figures show the average normalized root mean squared testing error (NRMSE) of reservoir computers with their corresponding standard deviations as we vary either the number of input frequencies $K$ or the number of nodes NN. Mean and standard deviations are computed from 20 independent runs. The yellow curve represents the NRMSE of the linear benchmark, the orange curve corresponds to the nonlinear RC with ReLU activation, the blue curve shows the nonlinear RC with tanh activation, and the purple curve indicates the optimized LRC. a) The number of nodes is fixed at $N=50$ while the number of input frequencies varies. b) The number of input frequencies is fixed at $K=10$ while the number of reservoir nodes varies. The duration of the testing phase is taken to be one third of the duration of the training phase.
Figure (5)(a) shows that the lowest testing error is achieved either by the tanh-based nonlinear RC in the case of a small number of frequencies $K$ (blue curve) or by the optimized LRC for larger values of $K$ (purple curve). The linear RC benchmark (yellow curve) and the nonlinear RC with ReLU activation (orange curve) appear to underperform the other two methods over the entire range of the frequencies we have tested. Figure (5)(b) shows that the optimized LRC (purple curve) achieves the lowest testing error, whereas the tanh-based nonlinear RC (blue curve) shows a greater reduction in error as the number of nodes increases. On the other hand, the linear RC benchmark (yellow curve) and the nonlinear RC with ReLU activation (orange curve) both appear to underperform the other two methods over the entire range of $N$ we have tested. An important observation from Fig. (5)(b) is that the optimized LRC achieves significantly lower NRMSE with far fewer nodes compared to the tanh-based RC. This reduction in required network size directly makes the proposed optimized LRC more efficient for real-time and resource-constrained applications.
Standard RCs with a tanh or ReLU reservoir are computationally fast to train, but their performance is limited by the randomly initialized reservoir. In contrast, the optimized linear RC entails a substantially higher computational cost during the optimization phase, as the underlying problem is nonconvex and highly sensitive to the initial guess (see Table 6 in the Supplementary Material). Achieving a near-optimal solution therefore requires evaluating multiple initializations, with the number of required guesses increasing with the network size, which explains the long runtimes observed. For a given task, however, this optimization is performed only once; once the reservoir has been optimized, the associated computational cost does not need to be incurred again.
Despite this computational cost, the optimized RC achieves lower estimation errors and attains a given level of performance with fewer nodes (reducing memory and deployment runtime). Since the optimization is performed offline, the resulting reservoir can be used efficiently in real-time or resource-constrained applications.
@@ 3.3 Effect of the Cost Function Components on the Optimization Performance
The optimization formulation, as shown in Eq. (25), consists of three terms. The second and third terms have associated weights $\beta_{1}\geq 0$ and $\beta_{2}\geq 0$, respectively. Therefore, it is important to analyze the impact of these coefficients on the overall optimization process. Figure (6) (a) demonstrates the effect of the second and third terms’ weights on the optimization of the error. Figure 6 (a), shows that reducing the weight of the second term ($\beta_{1}$) leads to a decrease in optimization error. This suggests that allowing larger output weights enables the reservoir to more effectively learn the signal, thereby improving performance. Additionally, increasing the weight of the third term ($\beta_{2}$) also leads to a reduction in error. This is because the third term, a heuristic algebraic term, encourages the optimization to produce distinct eigenvalues, which is a desirable property. We see that $\beta_{1}$ has a more substantial impact on optimization. Additionally, considering the heuristic role of $\beta_{2}$, its contribution should only be effective if it is nonzero and not too large. This balance can be seen in the reduction in error near the bottom right of Figure (6) (a).
@@ 3.4 Effect of $\lambda$ on Reservoir Performance
Performing a sensitivity analysis of the reservoir’s eigenvalues is important for assessing both robustness and generalization. High sensitivity means that small perturbations in eigenvalues can lead to large errors, making the system fragile. It may also indicate overfitting, where the reservoir performs well only under narrowly tuned conditions. Evaluating sensitivity helps ensure stable and reliable performance. Since the optimization formulation focuses on minimizing the error by tuning the reservoir’s eigenvalues and output weights, it is essential to understand how sensitive the reservoir computer is to the perturbations applied to the optimal eigenvalues. In order to perform the sensitivity analysis, a small perturbation is added to each one of the eigenvalues,
$\lambda_{perturbed}=\lambda_{opt}+\rho\delta,$ (26)
where $\delta$ is a number randomly chosen from a Normal distribution and $\rho\geq 0$ is a tunable parameter.
Figure 6: a) Average reservoir training NRMSE as a function of $\beta_{1}$ and $\beta_{2}$ for an $N=10$-node reservoir with $\gamma=5.5$. Each plotted data point, corresponding to a different selection of $\beta_{1}$ and $\beta_{2}$, is an average over 10 different runs. b) Sensitivity of the reservoir training NRMSE to perturbations applied to the optimal eigenvalues, based on Eq. (26). Each plotted data point is an average over 10 different runs.
Figure (6) (b) indicates that the optimal LRC training NRMSE remains largely unaffected by small perturbations applied to the optimal eigenvalues, ie., for $\rho$ that does not exceed a threshold value. Only significant deviations from these optimal values cause noticeable increases in the NRMSE. This behavior suggests that the reservoir is robust and stable around the optimized eigenvalues, making it resilient to minor noise or implementation inaccuracies. Consequently, the optimized reservoir parameters provide a reliable solution, ensuring consistent performance even when slight variations occur in practice.
3.5 Effect of the Number of the Training Steps $T$ on Reservoir Performance
A notable advantage of reservoir computing is its ability to learn from signals of short duration, which reduces the required data length and decreases the overall computational effort during training. Consequently, it is important to evaluate how the duration of the training phase influences the performance of the optimized reservoir computer. To this end, we have assessed the performance of the optimized LRC using multiple numbers of training steps $T$, and the resulting performance metrics are reported in Table 2. The NRMSE values presented in the Table are averaged over 50 independent runs.
Training Steps Avg. Training NRMSE Avg. Test NRMSE
500 0.0071 0.0067
1000 0.0061 0.0059
2000 0.0059 0.0056
3000 0.0058 0.0054
Table 2: Average training and test NRMSE for different numbers of training steps. The NRMSE values presented in the Table are averaged over 50 independent runs.
From Table 2 we see that as the number of training steps $T$ decreases, the NRMSE increases slightly; however, this increase is not significant and can be considered negligible. This indicates that the proposed optimized RC is robust to variations in the duration of the training-phase, while maintaining accurate performance.
@@ 4 Conclusion
This work makes two main contributions. First, we demonstrate that the dynamics of a linear RC can be decoupled into a number of independent modes in a way that maintains the training error (and the testing error) unchanged. Second, we show that each of these modes can be individually optimized for a specific task when the decoupled RC is reformulated from the time domain into the frequency domain. This reformulation yields a much smaller readout matrix, since the number of input frequencies is typically much smaller than the number of time steps. In so doing, we show that the effect of the connectivity of the LRC on the training and testing error is completely captured by the eigenvalues of the reservoir adjacency matrix, which allows us to formulate an optimization problem in the eigenvalues only.
A common practice in reservoir computing is to tune the spectral radius by uniformly rescaling all the eigenvalues of the adjacency matrix jaeger2001echo. Our method can be seen as an extension of this technique, which involves fine-tuning each individual eigenvalue. Simulations on networks of varying sizes demonstrate that the optimized LRC significantly outperforms randomly constructed reservoirs in both the training and testing phases. Moreover, comparisons with established RC benchmarks indicate that the optimized reservoir not only achieves improved performance but also exhibits greater robustness, scalability, and modularity. A limitation of our approach is that the optimization of the eigenvalues is non-convex, making the results highly dependent on the choice of initial guesses and computationally costly. In this work, the initial guesses were chosen randomly, which increases computational time and can lead to sub-optimal results. Identifying a more suitable choice of the initial guess remains an open challenge for future work. Although not addressed here, our work can be generalized to the case of the next generation reservoir computing of Ref. Gauthier2021